Systems Staff Meeting Notes 3/30/98

Systems Staff Meeting Notes
March 30, 1998

Present: Rich Bennett, Brian Burton, Bill Covey, Destry Fudge, Debra Harris, Winston Harris, Sharon Henry, Bobbie Parker, Suzy Shaw

Announcements
  1. Sam Gowanís Retirement Party - Tuesday, March 31, 4:00 - 6:00 PM @ the Reitz Union.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Two replacement modules from CISCO have been received. The remainder should be here no later than Wednesday, April 1st.
  2. Printer problems - icons have been disappearing. Have to trash, then reinstall. Reported in West & Systems.
  3. Intruders & hackers on campus - has not occurred in the Libraries - only recently we received some threatening messages on public work stations in MSL & West. Some hackers appear to have military IDís .
  4. PKY SNA outage last Friday - it was a control unit problem, Lee handled it by phone.
  5. Preservation - Color LaserWriter problems - the humidity is too low in that area, causing gray & red streaks across the page. The color tints are malfunctioning. Service people will charge to come out if they find the humidity is too low. Will try to monitor the humidity, maybe move the unit to another location in Preservation if necessary.
  6. Fiber cables ordered for switch installation in West & Education.
  7. Office í97 licenses have been ordered - one per each individual work station. The Office Ď97 has had bugs in it - the Excel application has had recalculating problems. Bug fixes canít be applied to the server version, but will work on the individual.
Networking
  1. CORE - web cache put into production this weekend. Please watch for any odd effects.
  2. We are committed to the Cambridge search engine as well as First Search, which is statewide & negotiated until the end of June.
  3. NERDC has switched back to UUNet, problem was a fiber cut & has been fixed.
  4. LAC - digital phones - much improved over what we have.
  5. Preservation - spec to Barbara Oliver.
Software Projects Status
  1. Systems Locator - housed on the NT2 in Systems Files, Applications - can drag it to your desktop or use a pointer to insure you have an updated version.
  2. ILL migration to SQL - test run is set for next week, Winston has made some changes to the accounting tables today.
  3. OPS tracking - Carolyn, Winston, Tom, & Barbara Oliver all met & identified changes to be made.
  4. Brian, Winston & Carolyn met about Brianís software licensing database - will encapsulate the licenses then add features.
  5. Evelyn - inputting OPS data. Salary savings are being handled by SAMAS - will want to make sure they are committed.
  6. David Allerton - Mellon - not done yet.
  7. Calendar software - will drop, can be handled by Outlook. Systems personnel may want to check out its functions.
DP Coordinator
  1. Debra has mailed the three 9-track tapes & 3 disks to the patron at Stanford University. We are getting off the auditor sheet in calculating the cost. Joan will generate an invoice: 25% overhead fee + cost + staff time. Total cost to patron was $93.43.
  2. We are in the process of re-evaluating how auxiliary contracts run in the Library. Would like to be divorced from printing & copier contracts - handle them internally.
Calendar
  1. Lee out today.
  2. Winston out March 31 - April 1.
  3. Bobbie out April 2 - 3.
  4. Sharon out April 3 - 6.
  5. Debra out April 10.
  6. Winston & Rich out April 13 - 20.
  7. Brian out April 16 - 17.
  8. Sharon out April 24 - 27.
  9. Memorial Day observed May 25.
  10. Bobbie out June 8 - 19.
  11. Independence Day observed July 3.
  12. Debra out July 7 - 10.
  13. Bobbie out July 13 - 17.
  14. Brian out July 16 & 17.


Back to the Systems Homepage

Last updated March 7, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu